OpenStudy (anonymous):

Given the vectors v=<-3,4> and u=2i-4j Determine a)2v+u b)vdotu c)the angle between u and v.

OpenStudy (jdoe0001):

let's us for rewrite the 2nd vector, understanding that v = ai + bj = <a, b> so v=<-3,4> u = <2, -4>

OpenStudy (jdoe0001):

2v is just a scalar multiplication so 2 * <-3, 4> => < 2(-3), 2(4)>

OpenStudy (jdoe0001):

addition is just <a, b> + <c, d> => <a+c, b+d>

OpenStudy (jdoe0001):

the "dot product" <a, b> dot <c, d> => ac + bd <--- it's just a number

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

thanks so the first one would be for a i

OpenStudy (jdoe0001):

v=<-3,4> u = <2, -4> 2v < 2(-3), 2(4)> => <-6, 8> 2v+u <-6, 8 > + <2, -4> => <-6+2, 8+(-4)> = <-4, 4>

OpenStudy (jdoe0001):

v=<-3,4> u = <2, -4> now the "dot product" \(\large <-3, 4> \bullet <2, -4> \implies -3(2) + 4(-4) = -22\)

OpenStudy (jdoe0001):

now, to find the angle in between both vectors you'd need to use the formula $$ cos(\theta)=\frac{v\times u}{||v||\times ||u||} $$ where the "magnitude" of say "u" -> \(||u||=\sqrt{a^2+b^2}\)

OpenStudy (anonymous):

Ohhh ok i see thanks for your help

OpenStudy (jdoe0001):

$$ ||v|| = \sqrt{(-3)^2+4^2} \implies \sqrt{25} \implies 5\\ ||u||=\sqrt{2^2+(-4)^2} \implies \sqrt{20} \implies 4\sqrt{5} $$

OpenStudy (jdoe0001):

$$ ||v|| = \sqrt{(-3)^2+4^2} \implies \sqrt{25} \implies 5\\ ||u||=\sqrt{2^2+(-4)^2} \implies \sqrt{20} \implies 2\sqrt{5} $$ rather

OpenStudy (jdoe0001):

the product atop from the "dot product" of the numerator will give you a number the denominator multiplication of the magnitudes will give you another divide them, that will be the \(\large cos(\theta)\) to find \(\large \theta\) just get the \(cos^{-1}(\theta)\) to both left and right sides, and that'd be the angle